Lupo Corporation uses a job-order costing system with a single plantwide predetermined overhead rate based on machine-hours. The company based its predetermined overhead rate for the current year on the following data: Total machine-hours 30,000 Total fixed manufacturing overhead cost $ 252,000 Variable manufacturing overhead per machine-hour $ 2.10 Recently, Job T687 was completed with the following characteristics: Number of units in the job 10 Total machine-hours 30 Direct materials $ 675 Direct labor cost $ 1,050 The estimated total manufacturing overhead is closest to:

Multiple Choice

a. $315,000

b.$252,000

c. $252,002

d. $63,000

Boesenhofer, Inc., manufactures and sells two products: Product N6 and Product N7. The company has an activity-based costing system with the following activity cost pools, activity measures, and expected activity:

Estimated Expected Activity
Activity Cost Pools Activity Measures Overhead Cost Product N6 Product N7 Total
Labor-related DLHs $ 345,060 2,100 5,000 7,100
Machine setups setups 128,534 600 800 1,400
Order size MHs 107,598 4,100 3,800 7,900
$ 581,192



The activity rate for the Machine Setups activity cost pool is closest to:

a. 73.57 per setup

b. 19.81 per setup

c. 26.24 per setup

d. 28.32 per setup

Weirick, Inc., manufactures and sells two products: Product T8 and Product P4. The company has an activity-based costing system with the following activity cost pools, activity measures, and expected activity:

Estimated Expected Activity
Activity Cost Pools Activity Measures Overhead Cost Product T8 Product P4 Total
Labor-related DLHs $ 121,100 6,000 3,000 9,000
Production orders orders 57,110 700 500 1,200
Order size MHs 893,250 3,000 2,800 5,800
$ 1,071,460

The total overhead applied to Product P4 under activity-based costing is closest to: (Round your intermediate calculations to 2 decimal places.)

a. $836,140

b. $431,228

c. $495,403

d. $950,360

The Charade Corporation is preparing its Manufacturing Overhead budget for the fourth quarter of the year. The budgeted variable manufacturing overhead is $7 per direct labor-hour; the budgeted fixed manufacturing overhead is $77,000 per month, of which $15,200 is factory depreciation.

If the budgeted direct labor time for November is 7,200 hours, then the total budgeted manufacturing overhead for November is:

a. $142,600

b. $77,000

c. $112,200

d. $127,400
Answer #1

Lupo Corporation estimated total manufacturing overheads =

Fixed Manufacuring Overhead + ((Variable manufacturing overhead per machine hours) X Machine Hours)

$252,000 + (30,000 X $2.10) = $315,000

Thus, the answer is option (a)

____________________________________________________________________________

The activity rate for the Machine Setups activity cost pool =

Total Machine Set ups Overhead / Total Set ups

=$128,534 / 1,400 = $91.81 per set up

*This seems answer (b) in multiple options in incorrectly published

________________________________________________________________________________

The total overhead applied to Product P4 under activity-based costing =

Labor related overheads= (121,100 / 9,000) * 3,000 = $40,366.67

Production orders overheads = (57,110 / 1200) * 500 = $23,795.83

Order Size overheads = (893,250 / 5800) * 2,800 = $431,224.14

The total overhead applied to Product P4 under activity-based costing = $495,386.64

The answer is (c) $495,403

_____________________________________________________________________________

total budgeted manufacturing overhead for November =

Fixed Manufacuring Overhead + ((Variable manufacturing overhead per direct labor hours) X Direct Labor Hours)

=$77,000 + (7,200 * 7)

=$127,400

The answer is (d) $127,400